Determine the new (1) quantity (2) inverse demand price and (3) inverse demand quantity if a
subsidy of $10 is paid on Chinese coal?

Suppose that demand and supply of Chinese coal are
Qd = 60 − 0.5P.
Qs = −2 + 2P.
1
Expert's answer
2020-11-10T10:34:29-0500

If a subsidy of $10 is paid on Chinese coal, then:

Q*s = -2 + 2(P - 10),

Q*s = Qd,

-2 + 2(P + 10) = 60 - 0.5P,

2.5P = 42,

P = $16.8 (2)

Q = 60 - 0.5×16.8 = 51.6 units (1 and 3).